How would you solve this?
The sum of two numbers is 42. Three times the frist number is the same as four times the other number. Find the numbers.

Answer by sofiyac(983)

let first number be x and second one y
x+y=42
3x=4y
solve the system of equations
x=42-y
3(42-y)=4y
126-3y=4y add 3y to each side
126=7y divide each side by 7
18=y
so if y is 18 then x=42-y so x=42-18=24